OpenStudy (anonymous):

Complete the square. x^2 - 12x

OpenStudy (anonymous):

To complete the square, first take half of the middle coefficient and then square it; add to both side.

OpenStudy (zzr0ck3r):

(x-6)^2 = 36

OpenStudy (anonymous):

Half of 12 is 6. 6² = 36

OpenStudy (zzr0ck3r):

you got the rest?

OpenStudy (zzr0ck3r):

they should teach this first, after you get good at it...its so much nicer

OpenStudy (anonymous):

can I just draw the answr choices and you tell me whats right?

OpenStudy (zzr0ck3r):

(x-6)^2 = 36 solve for x x-6 = +-6 x = +-6+6 x = 12 x = 0

OpenStudy (anonymous):